Founders Company has recently become aware of the large total discounts on its orders and would like to know the impact on profit. The company computed its operating profit as follows:

Net sales after discounts $200,000

Variable costs   80,000

Contribution margin $120,000  

Fixed costs 70,000

Operating profit 50,000

1) Suppose Founders could reduce its sales discounts to produce a 10% increase in net revenues but no changes in variable or fixed costs. By what percent would operating profits increase? How does the percentage compare to the percentage increase in net sales revenue?

2) Consider the ratio of operating profit to sales. How does this ratio relate to the percentage change in operating profit, for a given percentage change in the net sales revenue?
Answer #1

Current position

PARTICULAR. AMOUNT

sales. 2,00,000

Less

Variable cost. 80,000

Contribution. 1,20,000

Less

Fixed cost. 70,000

Profit. 50,000

REVISED POSITION - Increase in net revenue by 10%

​​​​​​PARTICULAR. AMOUNT

Sales ( 2,00,000*110%). 2,20,000

Less variable cost. 80,000

Contribution. 1,40,000

Less fixed cost 70,000

Profit. 70,000

NOTE- THERE IS NO CHANGE IN FIXED AND VARIABLE COST

THEREFORE INCREASE IN PROFIT BY RS. 20,000/. (70,000-50,000,)

% OF INCREASE IN EXISTING PROFIT 20,000/2,00,000=10%

% REVISED PROFIT 70,000/2,20,000=31.8%

2. RATIO :

EXISTING SCENARIO

OPERATING PROFIT / NET SALES = 50,000/2,00,000=1:4 OR 25%

REVISED SCENARIO

OPERATING PROFIT/NET SALES =70,000/2,20,000=7:22 OR 31.80%